### v1

# CoordField: Coordination Field for Agentic UAV Task Allocation In Low-altitude Urban Scenarios

Abstract
With the increasing demand for heterogeneous Unmanned Aerial Vehicle (UAV) swarms to perform complex tasks in urban environments, system design now faces major challenges, including efficient semantic understanding, flexible task planning, and the ability to dynamically adjust coordination strategies in response to evolving environmental conditions and continuously changing task requirements. To address the limitations of existing approaches, this paper proposes coordination field agentic system for coordinating heterogeneous UAV swarms in complex urban scenarios. In this system, large language models (LLMs) is responsible for interpreting high-level human instructions and converting them into executable commands for the UAV swarms, such as patrol and target tracking. Subsequently, a Coordination field mechanism is proposed to guide UAV motion and task selection, enabling decentralized and adaptive allocation of emergent tasks. A total of 50 rounds of comparative testing were conducted across different models in a 2D simulation space to evaluate their performance. Experimental results demonstrate that the proposed system achieves superior performance in terms of task coverage, response time, and adaptability to dynamic changes.
I Introduction
The growing demand for deploying Unmanned Aerial Vehicle (UAV) swarms to perform tasks such as pedestrian detection, vehicle tracking, and traffic signal monitoring in complex and dynamic urban environments has highlighted the advantages of multi-UAV systems over single UAV, particularly in terms of coverage, system redundancy, and operational efficiency. However, the increasing complexity of tasks and the highly dynamic nature of urban environments also pose significant challenges. Effective task allocation in such settings requires the systems to continuously reassign UAVs to new tasks, adapt to UAV failures and changing task priorities, and achieve decentralized intelligent coordination without relying on centralized control.
To address these challenges, various optimization algorithms have been developed for UAV swarm coordination, such as the Grey Wolf Optimizer [1] and the Whale Optimization Algorithm [2]. However, these methods are typically tailored to specific types of tasks and struggle to adapt to heterogeneous UAV clusters operating in highly dynamic environments. With the rise of large language models (LLMs), recent research has explored leveraging their powerful semantic understanding and tool-calling capabilities to address such problems. For example, several approaches attempt to utilize few-shot or zero-shot learning methods based on existing models, enabling LLMs to perform multi-round planning, itinerary calibration, and execution [3, 4]. Nevertheless, such methods still exhibit significant limitations when handling multiple tasks in complex and rapidly changing environments.
To address the above challenges, the emerging Agentic Artificial Intelligence(AI) offers a more flexible and unified approach [5]. This paradigm is driven by LLMs, and integrates core functionalities such as task decomposition, tool invocation, and multi-agent coordination and scheduling into a cohesive system [6, 7]. In this paper, we propose a coordination field multi-UAV agentic system architecture. The system leverages LLM to parse natural language instructions into executable tasks for UAV swarms, and employs a coordination field control strategy to achieve task-oriented autonomous navigation and collective coordination.
The main contributions of this work are as follows:
  • We propose CoordField, a coordination field-based UAV swarms task allocation method that employs continuously updated potential fields to represent both task urgency and UAV influence within urban environments. This approach enhances real-time responsiveness.
  • We present an agentic system for UAV swarms that enables high-precision task understanding from natural language descriptions and supports dynamic planning and deployment through specialized collaborative agents with distinct roles.
  • We evaluate our system against multiple baseline models under the same task scenarios, validating its superior coordination performance in urban environments.
II Preliminaries
II-A Agentic Systems
In recent years, Agentic AI has emerged as a novel intelligent paradigm characterized by autonomous perception, goal-driven planning, and adaptive decision-making, offering innovative and effective solutions for multi-agent collaborative tasks [8, 9]. Unlike traditional passive approaches, Agentic AI emphasizes continuous goal pursuit, contextual memory, adaptive tool use, and self-reflection capabilities of agents [10]. These abilities enable agents to operate with greater autonomy and coordination in dynamic environments. Agent systems powered by LLMs have demonstrated strong reasoning and planning capabilities, particularly in complex task decomposition and sequential execution [3]. These features provide a solid theoretical and methodological foundation for building autonomous multi-UAV systems. In complex urban task scenarios [11], Agentic AI can assist UAVs in understanding high-level instructions, distributing subtasks effectively, and rapidly reconfiguring coordination strategies in response to task changes or agent failures, thereby enabling truly autonomous collaboration.
II-B LLMs for Natural Language Task Parsing
With the rapid advancement of LLMs in natural language understanding and general reasoning, LLMs have demonstrated superior capabilities over traditional rule-based parsers in handling ambiguous semantics, conditional logic, and spatial language. They also exhibit strong generalization in task generation, even under limited domain-specific data. Researchers have developed several frameworks such as Code-as-Policies and Prompt2Action[12, 13], enabling LLMs to map natural language instructions into task-executable codes. These approaches often incorporate few-shot prompting and constraint prompting to improve the validity and safety of the generated outputs. Furthermore, the GSCE framework, which consists of Guidelines, Skill APIs, Constraints, and Examples, has demonstrated the practical value of prompt engineering in UAV task planning[14], guiding LLMs to generate physically feasible and constraint-compliant control programs by providing skill APIs, operational constraints, and illustrative examples.
II-C Field-Based Coordination and Allocation
Field-driven methods, particularly the Artificial Potential Field (APF) approach, have been widely applied in the field of mobile robotics[15]. In this approach, target points are modeled as attractive sources, obstacles as repulsive sources, and robots navigate by descending along the gradient of the resulting potential field, thereby enabling path planning and real-time obstacle avoidance. This approach has been applied applications in traffic flow modeling [16], where field-based methods are used to simulate the evolution of traffic dynamics. Vehicles are modeled as particles influenced by traffic density fields, velocity fields, and other factors, allowing the system to capture both macroscopic and microscopic variations more effectively [17].
Inspired by this line of thought, our work adopts a similar strategy by modeling task demand and UAV workload as continuous spatial fields. A dynamically evolving potential field is constructed based on the distribution of task urgency, guiding UAVs to move and allocate themselves through local gradient descent [18]. The advantage of this method lies in its decentralized structure, strong real-time adaptability, and high responsiveness to task changes, making it particularly well-suited for urban scenarios where tasks are unevenly distributed and dynamically changing. Compared with traditional discrete optimization approaches, field-driven methods offer a more flexible and scalable solution.
III Approach
In this work, we propose an Agentic system for UAV swarms, built upon a Coordination field mechanism is proposed to guide UAV motion and task selection, enabling decentralized and adaptive allocation of emergent tasks. methodology. The proposed agentic system consists of three primary modules, as illustrated in Fig. 1. It enables multi-layered agent coordination to interpret natural language inputs and perform task decomposition and allocation for UAV swarms in urban environments. In particular, to address the heterogeneity of UAV types and the complexity of urban traffic conditions, a field-driven strategy is adopted to guide different types of UAVs toward task regions that match their operational roles.
Figure 1: Coordination field Agentic System Implementation Process
III-A Semantic Understanding Module
The semantic understanding module is responsible for interpreting user-provided natural language inputs and converting them into structured task representations suitable for downstream planning and execution. Serving as the human–system interface, it enables intuitive interaction with the UAV swarm through language-based commands and ensures that high-level semantic information is accurately conveyed to subsequent modules in the system[19].
Upon receiving a user prompt, the LLM performs a task translation process, mapping vague or abstract natural language instructions into formalized constraints. Specifically, the LLM processes the input to extract key elements such as target location, task type, and priority level, and ultimately generates a structured task tuple in the following form:
$\text{Task}_{j}=(x_{j},y_{j},w_{j},\text{type}_{j})$
where $(x_{j},y_{j})$ denotes the spatial coordinates of the target, $w_{j}$ represents the task weight or priority, and $\text{type}_{j}$ refers to the semantic category of the task (e.g., patrol, tracking, inspection).
This formalization process acts as a bridge between high-level user intent and the low-level execution logic of the agentic system. It enables downstream modules, such as the planning and execution modules, to operate on a consistent and quantifiable task representation. Moreover, the module leverages the strong semantic understanding and reasoning capabilities of LLMs, laying a foundation for flexible task expansion and constraint integration.
III-B Planning Module and Coordination Field
The Planning Module is responsible for transforming the task tuples $(x_{j},y_{j},w_{j},\text{type}_{j})$ obtained from semantic parsing into a dynamic UAV coordination strategy. The Planning Agent first feeds the task tuples and UAV state data into the coordination field. This field-based framework consists of three submodules—perception mapping, task decomposition, and task assignment—which operate in a closed-loop manner to continuously adapt to evolving task demands and spatial distribution patterns. The results are then aggregated and delivered by the Planning Agent for execution.
Perception Mapping: The Planning Agent first reads relevant map information and UAV configurations, such as types, capabilities, and quantities. When a task emerges at any given time $t$, the Planning Agent groups related tasks and identifies a set of active tasks, each $Task_{j}$ located at position $(x_{j},y_{j})$ with a dynamic urgency weight $w_{j}(t)$.
To establish task density awareness in the environment, we construct a time-varying potential field $\phi(x,y,t)$ to represent the spatial intensity and distribution of high-priority task regions. This scalar field is composed of a sum of weighted Gaussian functions, defined as:
$\phi(x,y,t):=\sum_{j=1}^{M}w_{j}(t)\cdot\exp\left(-\frac{\left\|(x,y)-(x_{j}(t),y_{j}(t))\right\|^{2}}{2\sigma_{j}^{2}}\right)$
Here, $\sigma_{j}$ denotes the spatial influence radius of task $j$, reflecting the degree to which its urgency propagates across nearby areas. Notably, obstacles such as buildings are excluded from the field domain, and $\phi(x,y,t)$ is explicitly set to zero within these regions.
By constructing this continuous “task landscape”, UAVs are able to perceive and exploit the gradient of $\phi(x,y,t)$ to identify and navigate toward high-priority task areas.
Task decomposition: To convert the scalar potential field $\phi$ into actionable guidance for UAV navigation and control, we further construct a time-varying vector field $\mathbf{v}(x,y,t)$, which represents the motion velocity field of the UAV swarm. During the process, the following motion equation is defined based on the Navier–Stokes equations in fluid dynamics:
$\frac{d\mathbf{v}}{dt}:=-\frac{1}{\rho}\nabla p+\nu\nabla^{2}\mathbf{v}+\mathbf{F}_{\text{task}}(\phi)$
In this model, $\mathbf{v}(x,y,t)=(v_{x},v_{y})$ denotes the velocity vector at each point in the 2D space, indicating the direction and magnitude that a UAV should follow at a given position. The parameter $\nu$ represents the fluid viscosity, which governs the smoothness and responsiveness of the velocity field: higher viscosity results in smoother flows but slower response. $\mathbf{F}_{\text{task}}$ is the external force term used to steer UAVs toward regions with higher $\phi$ values, and it is defined as:
$\mathbf{F}_{\text{task}}=k\nabla\phi$
where $k$ is a scaling coefficient that adjusts the attraction strength. $\rho$ is the nominal fluid density, which can be normalized in this system for simplification.
By constructing such a velocity field, UAVs do not need to explicitly plan trajectories. Instead, they can simply follow the local flow direction, naturally converging toward high-demand task regions. Meanwhile, the inclusion of the diffusion term $\nu\nabla^{2}\mathbf{v}$ helps prevent over-concentration of UAVs in a single area, thereby enhancing spatial distribution and system robustness.
Task Assignment: To further enhance coordination efficiency among multiple agents and prevent resource redundancy or execution conflicts caused by UAVs clustering in the same task region, we introduce a local vortex mechanism around each UAV to serve as a dynamic repulsion control strategy. Specifically, each UAV $i$ generates a rotational field centered at its current position $(x_{i},y_{i})$, defined as:
$\omega_{i}(r):=\frac{\Gamma_{i}}{2\pi r}\exp\left(-\left(\frac{r}{r_{0}}\right)^{2}\right)$
where $r$ is the radial distance from UAV $i$, and $r_{0}$ denotes the influence radius of the field. $\Gamma_{i}$ is the circulation strength generated by the UAV, which depends on its capability score $c_{i}(t)$ and the local potential value $\phi(x_{i},y_{i},t)$, calculated as:
$\Gamma_{i}(t):=\frac{c_{i}(t)\cdot\phi(x_{i}(t),y_{i}(t),t)}{\sum_{j=1}^{N}c_{j}(t)}$
Based on this field, the system computes a tangential velocity component $v_{\theta,i}(r)$, which forms a repulsive velocity around each UAV and models the inter-agent avoidance dynamics:
$v_{\theta,i}(r):=\frac{\Gamma_{i}}{2\pi r}\left(1-\exp\left(-\left(\frac{r}{r_{0}}\right)^{2}\right)\right)$
Finally, the control velocity vector of each UAV is obtained by combining the global guiding velocity field $\mathbf{v}(x,y,t)$ with the superimposed tangential repulsive velocities generated by all UAVs:
$v_{\text{new}}(x,y,t)=v(x,y,t)+\sum_{i=1}^{n}v_{\theta,i}(x,y,t)$
This mechanism ensures that UAVs are not only guided toward high-priority task regions but also mutually repelled through localized vortex effects, thereby enabling dynamically stable task allocation and coordination behavior in a fully decentralized setting. The system thus achieves high adaptability and strong scalability.
III-C Execution Module
The core task of the Execution Module is to translate the outputs of the planning module, including the flow field vector $\mathbf{v}_{\text{new}}(x,y,t)$ and the UAV task assignment information, into real-time control commands executable by either physical or simulated UAVs. This process is managed by the Execution Agent, which communicates with the underlying control system through standardized API interfaces.
During the task execution phase, each UAV samples a local velocity vector from the precomputed control field $\mathbf{v}_{\text{new}}(x,y,t)$ based on its current position $(x,y,t)$, and uses this vector as its next navigation direction. The Execution Agent then generates flight control commands accordingly and dispatches them to the UAV platform (either physical or virtual) via middleware such as MAVSDK, ROS2, or AirSim. These control commands can be formalized as:
$\text{UAV}(x,y,t)\rightarrow\text{Map}(x^{\prime},y^{\prime},t^{\prime})$
which maps the current state to a new target position and time step.
Based on the task type $\text{type}_{j}$, UAVs are assigned distinct execution modes, such as performing area patrol or tracking specific targets. These operational differences are embedded in the command structure during instruction generation and are reflected in the API parameters.
To maintain adaptability in dynamic environments, each UAV continuously reports its status, including position, velocity, and execution state, to the control system. These updates are used to refresh the task potential field $\phi(x,y,t)$ and the flow field vector $\mathbf{v}(x,y,t)$ in real time. When the environment changes, such as when new task targets appear, current tasks are completed, or obstacles are introduced, the system recalculates the relevant fields and delivers updated control instructions through the Execution Agent.
This closed-loop perception–planning–execution cycle operates at a high frequency, ensuring the system exhibits strong responsiveness and decentralized dynamic coordination. Under this mechanism, UAVs are capable of re-planning, cooperative obstacle avoidance, and target adaptation in complex and ever-changing environments, thereby realizing truly adaptive intelligent behavior.
IV Experiments
Figure 2: 2D space simulation map.
To validate the proposed architecture, we conducted a series of experiments in a simulated urban environment. We first introduce the construction of the simulation environment, the UAV swarm configuration, the task generation process—including natural language instruction parsing—and the baseline methods used for comparison. We then describe the performance evaluation metrics and testing procedures. Finally, we present results from various experimental scenarios to demonstrate the system’s effectiveness in terms of task coverage, load balancing, and response latency.
IV-A Environment and UAV Configuration
The study is conducted within a custom-designed two-dimensional urban simulation environment, covering a $1000\times 1000$ grid space, as illustrated in Fig. 2. The environment features a structured city road network, with gray regions representing densely distributed buildings. Red dots mark traffic lights located at road intersections, while a large number of black points and blue rectangles denote dynamically moving pedestrians and vehicles, respectively. The roads are laid out in alignment with the city grid, creating a complex traffic and pedestrian flow environment for multi-agent interaction and coordination.
In this environment, UAVs operate within the airspace above the road network and maintain a consistent, safe flight altitude to simplify operations and ensure task execution uniformity. A total of 20 UAVs are deployed in the simulation, divided into two types: patrol UAVs and tracking UAVs, with 10 units each. While both types share identical flight performance characteristics, they differ in sensing and recognition capabilities. Patrol UAVs are better suited for wide-area scanning and anomaly detection tasks, whereas tracking UAVs are designed for precise following of specific targets. Each UAV is assigned a capability vector $\mathbf{c}_{i}(t)$, which dynamically represents its adaptability to different types of tasks.
To enhance the dynamism and complexity of the simulation, the environment is updated in real time with a time step $T$, allowing accurate modeling of pedestrian and vehicle movements, as well as changes in traffic signal states. Task instructions are issued via natural language input, such as “Please inspect the crowd and vehicles.” These instructions are parsed through the DeepSeek API, which interprets the natural language input and transforms it into structured task information, including parameters such as target location, task type, and priority.
IV-B Experimental Results and Analysis
Figure 3: Density map with Hotspots.
To validate the effectiveness of the proposed system, a series of systematic experimental evaluations are conducted, along with representative visualizations of key results. As shown in Fig. 3, the heatmap illustrates the task density distribution across the urban environment. Red and yellow regions indicate task hotspots, reflecting areas with a higher concentration of task demand, where the system dynamically reallocates UAV resources accordingly.
Figure 4: UAV task assignment trajectory route.
Fig. 4 presents the optimized UAV trajectories. Each blue trajectory connects a starting point (green triangle) to a task target (red circle), demonstrating the effect of global path optimization based on minimum travel distance. The results show that the UAVs efficiently cover multiple task points while maintaining a reasonable spatial distribution, thereby avoiding congestion and conflict.
The experimental evaluation metrics include the following 4 aspects:
1) Task Parsing Accuracy (TPA): This metric measures the proportion of natural language instructions that are correctly parsed into structured task representations. It reflects the system’s ability to interpret high-level human commands, which is critical for enabling intuitive human–agent interaction and ensuring accurate downstream execution.
2) Coverage Efficiency (CE): This quantifies the degree to which UAVs effectively cover the areas defined by the task potential field $\phi(x,y,t)$. It is computed as the ratio between the UAV-covered region and the total task-defined region. High coverage efficiency indicates the system’s effectiveness in spatial distribution and responsiveness to task urgency.
3) Task Load Balance (TLB): This evaluates how evenly the task load is distributed across all UAVs, using the norm of the task allocation density gradient $\nabla\rho$. A balanced load prevents overutilization of individual UAVs, reduces task redundancy, and ensures system robustness in dynamic environments.
4) UAV Utilization Rate (UUR): This metric represents the proportion of UAVs actively engaged in task execution relative to the total available fleet. Higher utilization reflects efficient agent coordination and resource deployment, which is essential for maximizing mission effectiveness and operational economy.
TABLE I: Comparison of Task Performance Metrics Across Models
Model TPA CE TLB UUR
Our method 96% 95% 0.8 97%
Deepseek-v3 70% 78% 2.1 85%
GPT-4o 74% 85% 1.5 82%
Claude-3-7-Sonnet 76% 80% 1.6 84%
LLaMA-4-Scout 72% 79% 1.9 83%
Gemini-2.5-Pro 73% 81% 1.7 84%
  • Model versions: Deepseek-v3 (25.03.24), GPT-4o (24.05.13), Claude-3-7-Sonnet (25.02.19), LLaMA-4-Scout (25.04.05), Gemini-2.5-Pro-Preview (25.03.25).
The experimental evaluation on 50 natural language command tasks demonstrates that the proposed coordination field agentic system outperforms all baseline models across multiple key performance metrics as shown in Table 1. Specifically, it achieves the highest TPA of 96%, significantly exceeding GPT-4o (74%)[20], Claude-3-7-Sonnet (76%)[21], and Deepseek-v3 (70%)[22], indicating its superior semantic understanding and robustness in converting unstructured commands into executable UAV instructions. In terms of CE, the system maintains an average rate of 95%, consistently outperforming other models such as GPT-4o (85%) and LLaMA-4-Scout (79%)[23], and remains highly reliable even under task-burst scenarios with over 80% effective coverage. With respect to TLB, our method demonstrates the lowest task distribution deviation of only 0.8, compared to 2.1 and 1.6 observed in Deepseek-v3 and Claude-3-7-Sonnet respectively, which significantly reduces resource clustering and enhances spatial coordination. Additionally, the UUR reaches 97%, confirming that the majority of UAVs remain actively engaged in task execution, whereas competing models such as GPT-4o and LLaMA-4-Scout fall below 85%. These results collectively validate the system’s strong coordination capability, dynamic adaptability, and overall efficiency in managing heterogeneous UAV swarms within complex urban environments.
V conclusion
This paper proposes a coordination field Agentic system for UAV task coordination. The system demonstrates efficiency, stability, and adaptability, enabling real-time parsing of natural language instructions and efficient task allocation via potential fields, fluid-based velocity fields, and vortex-based guidance. Experimental results verify its superior performance in complex and dynamic urban environments. In future work, we plan to extend the system to three-dimensional scenarios and real-world UAV swarms.